Say that you were to cross a few plants from the F1 generation that you got after the crossing true breeding PP with true breeding pp (PP * pp the true breeding purple plants (PP) with the true breeding white plants (pp) )and end up with an F2-generation of 437 purple-flowered individuals and 163 white-flowered individuals. Perform a Chi Square test on your data to see if there is a significant difference between your expected F2-generation and your actual F2-generation.

Observed Expected (Observed minus expected) (Observed-expected)^2 (Observed-expected)^2 / expected
Purple Flowers
White Flowers
X^2=

What would the p-value be here and would it be significant or insignificant also would you reject/accept it and why? Please explain as well with numerical data please if possible!
Answer #1

Total number of individual are 437+163= 600

Expected white(1/4) ×600= 150

Expected purple = (3/4) ×600= 450

Now look in the image for chi square test:

Now that chi square value from table at degree of freedom =n-1 = 2-1 = 1 and p=0.05 is 3.841

Here n= number of phenotype = 2

As our chi square experimental value is less than chi square value from table. That's why we accept our results or hypothesis. It means there is not found significant difference between observed and expected numbers.